OpenStudy (anonymous):

Can someone explain how to find the constant term using the binomial theorem?

OpenStudy (anonymous):

\[(6x + \frac{ 1 }{ 3x })^6\]

OpenStudy (anonymous):

\[(a+b)^n=\sum_{k=0}^n\begin{pmatrix}n\\k\end{pmatrix}a^{n-k}b^k\] Here, \(a=6x\), \(b=\dfrac{1}{3x}\), and \(n=6\). The constant term will occur when the powers of \((6x)^{6-k}\) and \(\left(\dfrac{1}{3x}\right)^k=(3x)^{-k}\) cancel out. This happens when \(6-k+(-k)=0~\Rightarrow~k=3\). So, you'd plug in \(k=3\) into the summation formula: \[\begin{pmatrix}6\\3\end{pmatrix}(6x)^{6-3}\left(\frac{1}{3x}\right)^3=\cdots\]

OpenStudy (anonymous):

Uhh, I don't understand what this:  symbol means

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

It's a summation. For example, \[\sum_{n=0}^2n=0+1+2\] Have you been formally introduced to the binomial theorem?

OpenStudy (anonymous):

Yes, with combinations, where (x+y)^n = nC0x^n + nC1x^(n-1)y, +...... nCny^n

OpenStudy (anonymous):

Ah, well my summation notation is a way of saying the same thing. \[\sum_{k=0}^n {}_nC_{k}x^{n-k}y^k={}_nC_0x^n+{}_nC_1x^{n-1}y+\cdots+{}_nC_ny^n\] \(k\) is called an index, where \(k=0,1,2,3,...\)

OpenStudy (anonymous):

Do you understand how to find the constant term though?

OpenStudy (anonymous):

Sort of

OpenStudy (anonymous):

Let me try to do another example

OpenStudy (anonymous):

\[(x^2 + \frac{ 1 }{ x })^4\]

OpenStudy (anonymous):

a = x^2 b = 1/x n = 4

OpenStudy (anonymous):

and k would be found by doing 4-k + -k = 0 = -2k = 4 = k = -2?

OpenStudy (anonymous):

\(k\) must be positive! \[4-k+(-k)=0~\Rightarrow~4=2k~\Rightarrow~k=2\]

OpenStudy (anonymous):

Whoops, my mistake, I forgot to flip the sign when I moved it over.

OpenStudy (anonymous):

Also, I got 160 for the first example? Is that correct?

OpenStudy (anonymous):

So, if \(k=2\), then the constant term will be \[{}_4C_2\left(x^2\right)^{4-2}\left(\frac{1}{x}\right)^2=\cdots\] hmm, looks like you made a mistake here. This gives you a nonconstant term. Are you supposed to find a constant term? And yes, you're right about 160.

OpenStudy (anonymous):

Here is a screenshot of the question http://gyazo.com/0ae4153301d80ff9a17baf1d5c71ca5e

OpenStudy (anonymous):

Okay, the \(x^2\) must be tripping you up. From the formula, you have \[\left(x^2+\frac{1}{x}\right)^4={}_4C_0(x^2)^4+{}_4C_1(x^2)^3\left(\frac{1}{x}\right)+{}_4C_2(x^2)^2\left(\frac{1}{x}\right)^2\\ ~~~~~~~~~~~~~~~~~~~~~~~+{}_4C_3x^2\left(\frac{1}{x}\right)^3+{}_4C_4\left(\frac{1}{x}\right)^4\] Each term will contain a power of \(x\), so there is not constant term. Well, there is, but it's 0.

OpenStudy (anonymous):

Each term will be of the form \[{}_4C_k\left(x^2\right)^{4-k}\left(\frac{1}{x}\right)^k\] Rewriting a bit, you have \[{}_4C_k\left(x^2\right)^{4-k}\left(x^{-1}\right)^k\\ {}_4C_k\left(x\right)^{2(4-k)}\left(x\right)^{-k}\\ {}_4C_k\left(x\right)^{2(4-k)-k}\] The constant term would appear when the exponent is 0: \[2(4-k)-k=0\\ 8-3k=0\\ 8=3k\] Although there is a solution for \(k\) here, keep in mind that \(k\) must be a whole number (\(k=0,1,2,3,...\)). So, the constant term is zero.

OpenStudy (anonymous):

Okay, I want to try one more problem to make sure I understand.

OpenStudy (anonymous):

\[\left( x^2 + \frac{ 1 }{ x } \right)^6\]

OpenStudy (anonymous):

For this one, you can use the same reasoning as the previous problem. The only difference is that \(n=6\), and not 4.

OpenStudy (anonymous):

6Ck (x^2)^(6-k) * (1/x)^k

OpenStudy (anonymous):

2(6-k) - k = 12-2k-k 12 = 3k k = 4

OpenStudy (anonymous):

Yes, so now you just plug in \(k=4\) into \( {}_6C_k\left(x\right)^{2(6-k)-k}\)

OpenStudy (anonymous):

I got 15, is that correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thank a ton, there are quite a bit of questions like this, and you've help me a lot

OpenStudy (anonymous):

You're welcome!